# QOSST: A Highly-Modular Open Source Platform for Experimental Continuous-Variable Quantum Key Distribution

### v1

@@ III.5 Bob
The beat frequency is the frequency difference between the two lasers, and can typically move by a few $\mathrm{MHz}\text{/}$ if the lasers are not very stable. This was not the case in our implementation thanks to the stability of our lasers. The sequence is unshifted by $f_{\mathrm{beat}}$. This step is known as carrier frequency estimation. Then, the Zadoff-Chu sequence is found by cross-correlations, which gives a precise value for the beginning of the sequence. Knowing the number of symbols and the symbol rate, the sequence is reduced to the useful sequence containing only the symbols. The rest of the DSP is performed by subframes because the change in $f_{\mathrm{beat}}$ can be relatively quick and we want to do the analysis on a time scale where $f_{\mathrm{beat}}$ can be considered constant. In particular, this subframe analysis is critical when the lasers used are not very stable. For each subframe, the frequencies of the two pilots are estimated again, and $f_{\mathrm{beat}}$ is also estimated again using Eq. (10). The first pilot is filtered and used as a phase reference, and a matching RRC filter is applied to the quantum data, after it has been shifted to baseband by $f_{\mathrm{beat}}+f_{\mathrm{shift}}$. Following this, the best sampling point is found by the maximal variance (eye diagram) method Kikuchi (2016). The phase is then corrected using the phase reference from the pilot (a uniform filter is applied to the phase reference for better performance). At this point, Alice’s and Bob’s symbols only differ by a global phase (and some noise) and the global phase correction is performed by testing all global rotations and keeping the one with the best covariance term $\langle XY\rangle$. More details on the DSP of both Alice and Bob can be found in appendix E.
@@
In this way, Bob ends up with $T$, $\xi$, $\langle n\rangle$, $\eta$ and $V_{el}$ and can proceed to the estimation of the secret key rate using Eq. (4) (using code from the skr package). We typically use the value $\beta=0.95$ for the reconciliation efficiency Jouguet et al. (2011).
@@
III.6 Secret Key Rate
The skr package provides classes with static methods to compute the secret key rate depending on the setup and the security assumptions. Three methods are currently available: untrusted homodyne detector Leverrier (2009), trusted homodyne detector Lodewyck et al. (2007) and trusted heterodyne detector Fossier et al. (2008), all in the asymptotic case. In particular, the code that is released in this package does not take into account discrete modulations and finite-size effects.
III.7 Simulations and Post-Processing
@@
IV Results
To assess the performance of CV-QKD driven by QOSST, we tested and benchmarked it using the experimental setup presented in section II. For this, we used five channel configurations. In four of them, the PM VOA emulates an optical fiber of 00 0, 55 55, 1010 1010 and $25\;$\displaystyle\mathrm{km}\text{/}$$ at $0.2\;$\displaystyle\mathrm{dB}\text{/}\mathrm{km}$$ (the actual attenuation values were in practice slightly higher). The last configuration uses a fiber spool of $25\;$\displaystyle\mathrm{km}\text{/}$$ (total attenuation: $5.22\;$\displaystyle\mathrm{dB}\text{/}$$) and a manual polarisation controller to compensate for the polarisation transformation of the fiber.
@@ IV Results
For the roll-off factor, we see that a value below 0.20.2 0.20.2 induces a slightly higher excess noise, and for high values a steep increase of the excess noise is observed. We can explain this effect by recalling that the bandwidth of the signal after the RRC filter is given by Eq. (8). Since in this experiment the frequencies of the pilots are fixed, Eq. (7) is violated when $\beta_{\mathrm{RRC}}$ becomes too big. In practice, the quantum data and the pilots start to be superposed in frequency, increasing the noise on the quantum data and reducing the precision of the phase compensation algorithm. This effect starts when $\beta_{\mathrm{RRC}}=0.65$, which corresponds to having $f_{\mathrm{shift}}+B_{s}/2=182.5\;$\displaystyle\mathrm{MHz}\text{/}$$ and is the first roll-off value where $f_{\mathrm{shift}}+B_{s}/2>f_{\mathrm{pilot},1}=180\;$\displaystyle\mathrm{MHz}\text{/}$$.
@@
Finally for the pilot amplitudes, a too low value results in a very high excess noise. This is due to the fact that either the frequency of the pilots cannot be properly recovered, or the SNR of the pilots is too low to get a good phase compensation. Also, the excess noise increases with the pilot amplitudes, which can be explained by frequency cross-talk between the quantum data and the pilots. The optimal value depends mainly on the SNR at the receiver, which means that this value should be optimized for every distance. This optimization was done for each of the 5 configurations. This first coarse optimisation was also followed by a fine optimisation (in this example in the region 0.05-0.15). The amplitude of the pilots is given in arbitrary units, and the actual amplitude of the pilot is proportional to this value. The optimal value was found to correspond to $\sim 12\;$\displaystyle\mathrm{dB}\text{/}$$.
@@
$\beta_{\mathrm{RRC}}$ 0.50.5 0.50.5
$R_{s}$ $100\;$\displaystyle\mathrm{MBaud}\text{/}$$
$f_{\mathrm{shift}}$ $100\;$\displaystyle\mathrm{MHz}\text{/}$$
$f_{\mathrm{pilot},1}$ $180\;$\displaystyle\mathrm{MHz}\text{/}$$
$f_{\mathrm{pilot},2}$ $200\;$\displaystyle\mathrm{MHz}\text{/}$$
$L_{ZC}$ 39893989 39893989
$R_{ZC}$ 55 55
Acquisition time $50\;$\displaystyle\mathrm{ms}\text{/}$$
Shot noise time $10\;$\displaystyle\mathrm{ms}\text{/}$$
DAC rate $2\;$\displaystyle\mathrm{GSa}\text{/}\mathrm{s}$$
ADC rate $2.5\;$\displaystyle\mathrm{GSa}\text{/}\mathrm{s}$$
@@
In Fig. 6, we plot the excess noise for the experiments using the VOA with attenuation corresponding to $25\;$\displaystyle\mathrm{km}\text{/}$$ and the fiber spool. The estimation of the excess noise is rather stable. The stability of the estimation of $\langle n\rangle$, $T$, shot noise and $V_{el}$ were also checked. The results in terms of the secret key rate are given in Fig. 7.
Figure 6: Excess noise at Bob’s side for the experiments with the VOA (equivalent to 25 km) and the 25 km fiber spool. Top: experiment with the VOA. One frame where the DSP failed was removed from the 100 frames. The average excess noise at Bob’s side is $0.0062\;$\displaystyle\mathrm{SNU}\text{/}$$. Bottom: experiment with the fiber. Only the first 50 frames where the experiment was stable are shown. During this time, the average excess noise at Bob’s side is $0.0072\;$\displaystyle\mathrm{SNU}\text{/}$$.
Experiment $\bm{\xi_{B}}$ (SNU) $\bm{K_{\infty}}$ (MBit/s) $\bm{K_{\mathrm{FSE}}}$ (MBit/s)
VOA $0\;$\displaystyle\mathrm{km}\text{/}$$ 0.00950.0095 0.00950.0095 22.422.4 22.422.4 17.717.7 17.717.7
VOA $5\;$\displaystyle\mathrm{km}\text{/}$$ 0.00910.0091 0.00910.0091 11.911.9 11.911.9 5.825.82 5.825.82
VOA $10\;$\displaystyle\mathrm{km}\text{/}$$ 0.00760.0076 0.00760.0076 6.356.35 6.356.35 2.552.55 2.552.55
VOA $25\;$\displaystyle\mathrm{km}\text{/}$$ 0.00620.0062 0.00620.0062 1.431.43 1.431.43 00 0
Fiber $25\;$\displaystyle\mathrm{km}\text{/}$$ 0.00720.0072 0.00720.0072 1.171.17 1.171.17 00 0
@@
Figure 7: Achievable secret key rate in the asymptotic and finite-size regime. The security parameter is $\varepsilon=10^{-10}$. For each experiment, the actual attenuation performed by the VOA was always slightly above the targeted one. For the experiments with the VOA, the extra attenuation corresponds to a slightly higher attenuation coefficient on the actual fiber, and a small mismatch in polarisation that is not included in the efficiency. For the experiment with the fiber spool and the VOA at $25\;$\displaystyle\mathrm{km}\text{/}$$ the finite-size secret key rate was zero.
The average results are summarized in table 2. The secret key rates are given for both the asymptotic and finite-size case. For the experiments with the fiber spool and the VOA at $25\;$\displaystyle\mathrm{km}\text{/}$$, the finite-size key rate is zero using the worst-case estimators. However, it is also possible to measure the finite-size key rate for each frame using the framework from Leverrier et al. (2010), similarly to how the system would operate in an operational setting, but once again it is not possible to extract a secret key rate for the number of symbols used here. We obtain a positive average when considering at least $10^{7}$ symbols (average finite-size key rate in order of hundreds of $\mathrm{kbit}\text{/}\mathrm{s}$), but it would be challenging to reach this block size over the same number of frames, since it would roughly multiply the DSP, and hence the experiment time, by 10. Increasing the block size will require to perform optimisation on the overall software in order to reduce the computation time (and memory usage).
V Discussion
@@
Unfortunately the experiments at $25\;$\displaystyle\mathrm{km}\text{/}$$ do not not give any key rate in the finite-size case with the worst-case estimator method. As was briefly mentioned above, another method is to follow the analysis in Leverrier et al. (2010) for each frame. For $N=10^{6}$, which was the number of symbols in our frame, we unfortunately also got a key rate of zero for each frame. However, increasing $N$ for the analysis, we start to get some frames with a positive finite-size key rate. For the experiment with 25 km, 40 frames out of 100 at $N=10^{7}$ and then all the frames at $N=10^{8},10^{9},10^{10}$. For the experiment with the fiber, there were respectively 14, 77, 93 and 94 frames with a positive finite-size key rate at $N=10^{7},10^{8},10^{9},10^{10}$. We now give the average key rate over all the frames (and in parentheses the average over the positive frames) for the VOA experiment at 25 km: $96\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$ ($240\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$), $817\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$, $1.1\;$\displaystyle\mathrm{Mbit}\text{/}\mathrm{s}$$, $1.2\;$\displaystyle\mathrm{Mbit}\text{/}\mathrm{s}$$ for $N=10^{7},10^{8},10^{9},10^{10}$ and for the experiment with the fiber: $39\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$ ($280\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$), $475\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$ ($617\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$), $718\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$ ($772\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$)$, $803\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$ ($854\;$\displaystyle\mathrm{kbit}\text{/}\mathrm{s}$$) with the same $N$. This makes us confident that one way to get better results would be to increase the number of points that are used for the estimation of $T$ and $\xi$. We indeed checked that the variance of those two variables has a dependence over the number of samples used to estimate them. This would reduce the confidence interval and give better worst-case estimators and subsequently better key rates.
@@
One of the current limitations of the QOSST software is the achievable distance. In practice the DSP can work at least until $10\;$\displaystyle\mathrm{dB}\text{/}$$ but at that value the estimation of the distance is imprecise and the excess noise too high to get a secret key. However, the distance could be improved either by improving the DSP or by changing the generation of the pilots. Indeed, we believe that one of the limitations comes from the finite amplitude difference that can be applied between the quantum symbols and the pilots. A solution could be to generate the pilots and the quantum symbols on different paths, or to send the pilots on the other polarisation, which would allow for more powerful pilots and less crosstalk, as suggested in Pi et al. (2023).
@@
The choice of the optimised parameters reported in this work was performed using a slightly older system with a DAC that was limited to $500\;$\displaystyle\mathrm{MSa}\text{/}\mathrm{s}$$ (Keysight M330A) and with a bandwidth of $200\;$\displaystyle\mathrm{MHz}\text{/}$$. This justifies the relatively low symbol rate used in the present work. The increase of the symbol rate (and hence the occupation bandwidth), allowed by the new DAC will be the subject of future investigation. We believe that there are no particular limitations coming from our software other than longer time and larger memory occupation for the DSP.
@@
Currently, the DSP (and the parameter estimation) is run for each frame before requesting the next frame, which means that two frames are temporally separated by the DSP time. We conducted an experiment on 20 frames to evaluate the DSP time and we measured two durations: the first is the DSP itself, which outputs the symbols for each subframe with a global phase difference, and the second is recovery of the global phase, which also requires Alice sending her data to Bob. The experiment is conducted with the same parameters as the experiment for $0\;$\displaystyle\mathrm{km}\text{/}$$ and requests $50\%$ of the symbols to Alice (500000500000 500000500000 symbols per frame). The average time for the DSP is $70\;$\displaystyle\mathrm{s}\text{/}\mathrm{frame}$$ and the average time for the symbol exchange and global phase correction is $164\;$\displaystyle\mathrm{s}\text{/}\mathrm{frame}$$ for a total time between frames of $234\;$\displaystyle\mathrm{s}\text{/}\mathrm{frame}$$. The second time can be reduced by exchanging less data for the parameter estimation. For example, taking $10\%$ of the data (i.e. 100000100000 100000100000 symbols for the parameter estimation), the DSP time is similar ($72\;$\displaystyle\mathrm{s}\text{/}\mathrm{frame}$$ on average) and the data exchange and global phase recovery has an average time of $101\;$\displaystyle\mathrm{s}\text{/}\mathrm{frame}$$, yielding a total of $173\;$\displaystyle\mathrm{s}\text{/}$$ between the frames. This has the advantage of reducing the time and increasing the key rate, but the estimators for the finite-size analysis are worse. Both times could be reduced by optimizing the code, or by using a different programming language.
VI Conclusion
We have presented a highly modular open source platform for experimental continuous-variable quantum key distribution that is able to exchange a secret key rate at metropolitan distances using the OSSB-GCMS-GG02 protocol with frequency multiplexed pilots and RF-heterodyne coherent detection. We obtain key rates of the order of the $\mathrm{Mbit}\text{/}\mathrm{s}$ for distances up to $25\;$\displaystyle\mathrm{km}\text{/}$$ at the asymptotic limit. Our platform can be easily extended to other CV-QKD protocols and hardware configurations, including with double polarisation and phase-diverse heterodyne detection. Our platform can be improved and extended by integrating techniques accessible in the near term.

### v3

@@ 3.5 Bob
The beat frequency is the frequency difference between the two lasers, and can typically move by a few $\mathrm{MHz}\text{/}$ if the lasers are not very stable. This was not the case in our implementation thanks to the stability of our lasers. The sequence is unshifted by $f_{\mathrm{beat}}$. This step is known as carrier frequency estimation. Then, the Zadoff-Chu sequence is found by cross-correlations, which gives a precise value for the beginning of the sequence. Knowing the number of symbols and the symbol rate, the sequence is reduced to the useful sequence containing only the symbols. The rest of the DSP is performed by subframes because the change in $f_{\mathrm{beat}}$ can be relatively quick and we want to do the analysis on a time scale where $f_{\mathrm{beat}}$ can be considered constant. In particular, this subframe analysis is critical when the lasers used are not very stable. For each subframe, the frequencies of the two pilots are estimated again, and $f_{\mathrm{beat}}$ is also estimated again using Eq. (10). The first pilot is filtered and used as a phase reference, and a matching RRC filter is applied to the quantum data, after it has been shifted to baseband by $f_{\mathrm{beat}}+f_{\mathrm{shift}}$. Following this, the best sampling point is found by the maximal variance (eye diagram) method [34]. The phase is then corrected using the phase reference from the pilot (a uniform filter is applied to the phase reference for better performance). At this point, Alice’s and Bob’s symbols only differ by a global phase (and some noise) and the global phase correction is performed by testing all global rotations and keeping the one with the best covariance term $\langle XY\rangle$. More details on the DSP of both Alice and Bob can be found in appendix E.
@@
In this way, Bob ends up with $T$, $\xi$, $\langle n\rangle$, $\eta$ and $V_{el}$ and can proceed to the estimation of the secret key rate using Eq. (4) (using code from the skr package). We typically use the value $\beta=0.95$ for the reconciliation efficiency [35].
@@
3.6 Secret Key Rate
The skr package provides classes with static methods to compute the secret key rate depending on the setup and the security assumptions. Three methods are currently available: untrusted homodyne detector [36], trusted homodyne detector [37] and trusted heterodyne detector [38], all in the asymptotic case. In particular, the code that is released in this package does not take into account discrete modulations and finite-size effects.
3.7 Simulations and Post-Processing
@@
4 Results
To assess the performance of CV-QKD driven by QOSST, we tested and benchmarked it using the experimental setup presented in section 2. For this, we used five channel configurations. In four of them, the PM VOA emulates an optical fiber of 00, $5$, $10$ and $25\;$\mathrm{km}\text{/}$$ at $0.2\;$\mathrm{dB}\text{/}\mathrm{km}$$ (the actual attenuation values were in practice slightly higher). The last configuration uses a fiber spool of $25\;$\mathrm{km}\text{/}$$ (total attenuation: $5.22\;$\mathrm{dB}\text{/}$$) and a manual polarisation controller to compensate for the polarisation transformation of the fiber.
@@ 4 Results
For the roll-off factor, we see that a value below $0.2$ induces a slightly higher excess noise, and for high values a steep increase of the excess noise is observed. We can explain this effect by recalling that the bandwidth of the signal after the RRC filter is given by Eq. (8). Since in this experiment the frequencies of the pilots are fixed, Eq. (7) is violated when $\beta_{\mathrm{RRC}}$ becomes too big. In practice, the quantum data and the pilots start to be superposed in frequency, increasing the noise on the quantum data and reducing the precision of the phase compensation algorithm. This effect starts when $\beta_{\mathrm{RRC}}=0.65$, which corresponds to having $f_{\mathrm{shift}}+B_{s}/2=182.5\;$\mathrm{MHz}\text{/}$$ and is the first roll-off value where $f_{\mathrm{shift}}+B_{s}/2>f_{\mathrm{pilot},1}=180\;$\mathrm{MHz}\text{/}$$.
@@
Finally for the pilot amplitudes, a too low value results in a very high excess noise. This is due to the fact that either the frequency of the pilots cannot be properly recovered, or the SNR of the pilots is too low to get a good phase compensation. Also, the excess noise increases with the pilot amplitudes, which can be explained by frequency cross-talk between the quantum data and the pilots. The optimal value depends mainly on the SNR at the receiver, which means that this value should be optimized for every distance. This optimization was done for each of the 5 configurations. This first coarse optimisation was also followed by a fine optimisation (in this example in the region 0.05-0.15). The amplitude of the pilots is given in arbitrary units, and the actual amplitude of the pilot is proportional to this value. The optimal value was found to correspond to $\sim 12\;$\mathrm{dB}\text{/}$$.
@@
$\beta_{\mathrm{RRC}}$ $0.5$
$R_{s}$ $100\;$\mathrm{MBaud}\text{/}$$
$f_{\mathrm{shift}}$ $100\;$\mathrm{MHz}\text{/}$$
$f_{\mathrm{pilot},1}$ $180\;$\mathrm{MHz}\text{/}$$
$f_{\mathrm{pilot},2}$ $200\;$\mathrm{MHz}\text{/}$$
$L_{ZC}$ $3989$
$R_{ZC}$ $5$
Acquisition time $50\;$\mathrm{ms}\text{/}$$
Shot noise time $10\;$\mathrm{ms}\text{/}$$
DAC rate $2\;$\mathrm{GSa}\text{/}\mathrm{s}$$
ADC rate $2.5\;$\mathrm{GSa}\text{/}\mathrm{s}$$
@@
In Fig. 6, we plot the excess noise for the experiments using the VOA with attenuation corresponding to $25\;$\mathrm{km}\text{/}$$ and the fiber spool. The estimation of the excess noise is rather stable. The stability of the estimation of $\langle n\rangle$, $T$, shot noise and $V_{el}$ were also checked. The results in terms of the secret key rate are given in Fig. 7.
Figure 6: Excess noise at Bob’s side for the experiments with the VOA (equivalent to 25 km) and the 25 km fiber spool. Top: experiment with the VOA. One frame where the DSP failed was removed from the 100 frames. The average excess noise at Bob’s side is $0.0062\;$\mathrm{SNU}\text{/}$$. Bottom: experiment with the fiber. Only the first 50 frames where the experiment was stable are shown. During this time, the average excess noise at Bob’s side is $0.0072\;$\mathrm{SNU}\text{/}$$.
Experiment $\bm{\xi_{B}}$ $\bm{K_{\infty}}$ $\bm{K_{\mathrm{FSE}}}$
(SNU) (MBit/s) (MBit/s)
VOA $0\;$\mathrm{km}\text{/}$$ $0.0095$ $22.4$ $17.7$
VOA $5\;$\mathrm{km}\text{/}$$ $0.0091$ $11.9$ $5.82$
VOA $10\;$\mathrm{km}\text{/}$$ $0.0076$ $6.35$ $2.55$
VOA $25\;$\mathrm{km}\text{/}$$ $0.0062$ $1.43$ 00
Fiber $25\;$\mathrm{km}\text{/}$$ $0.0072$ $1.17$ 00
@@
Figure 7: Achievable secret key rate in the asymptotic and finite-size regime. The security parameter is $\varepsilon=10^{-10}$. For each experiment, the actual attenuation performed by the VOA was always slightly above the targeted one. For the experiments with the VOA, the extra attenuation corresponds to a slightly higher attenuation coefficient on the actual fiber, and a small mismatch in polarisation that is not included in the efficiency. For the experiment with the fiber spool and the VOA at $25\;$\mathrm{km}\text{/}$$ the finite-size secret key rate was zero.
The average results are summarized in table 2. The secret key rates are given for both the asymptotic and finite-size case. For the experiments with the fiber spool and the VOA at $25\;$\mathrm{km}\text{/}$$, the finite-size key rate is zero using the worst-case estimators. However, it is also possible to measure the finite-size key rate for each frame using the framework from [20], similarly to how the system would operate in an operational setting, but once again it is not possible to extract a secret key rate for the number of symbols used here. We obtain a positive average when considering at least $10^{7}$ symbols (average finite-size key rate in order of hundreds of $\mathrm{kbit}\text{/}\mathrm{s}$), but it would be challenging to reach this block size over the same number of frames, since it would roughly multiply the DSP, and hence the experiment time, by 10. Increasing the block size will require to perform optimisation on the overall software in order to reduce the computation time (and memory usage).
5 Discussion
@@
Unfortunately the experiments at $25\;$\mathrm{km}\text{/}$$ do not not give any key rate in the finite-size case with the worst-case estimator method. As was briefly mentioned above, another method is to follow the analysis in [20] for each frame. For $N=10^{6}$, which was the number of symbols in our frame, we unfortunately also got a key rate of zero for each frame. However, increasing $N$ for the analysis, we start to get some frames with a positive finite-size key rate. For the experiment with 25 km, 40 frames out of 100 at $N=10^{7}$ and then all the frames at $N=10^{8},10^{9},10^{10}$. For the experiment with the fiber, there were respectively 14, 77, 93 and 94 frames with a positive finite-size key rate at $N=10^{7},10^{8},10^{9},10^{10}$. We now give the average key rate over all the frames (and in parentheses the average over the positive frames) for the VOA experiment at 25 km: $96\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($240\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $817\;$\mathrm{kbit}\text{/}\mathrm{s}$$, $1.1\;$\mathrm{Mbit}\text{/}\mathrm{s}$$, $1.2\;$\mathrm{Mbit}\text{/}\mathrm{s}$$ for $N=10^{7},10^{8},10^{9},10^{10}$ and for the experiment with the fiber: $39\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($280\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $475\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($617\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $718\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($772\;$\mathrm{kbit}\text{/}\mathrm{s}$)$, $803\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($854\;$\mathrm{kbit}\text{/}\mathrm{s}$$) with the same $N$. This makes us confident that one way to get better results would be to increase the number of points that are used for the estimation of $T$ and $\xi$. We indeed checked that the variance of those two variables has a dependence over the number of samples used to estimate them. This would reduce the confidence interval and give better worst-case estimators and subsequently better key rates.
@@
One of the current limitations of the QOSST software is the achievable distance. In practice the DSP can work at least until $10\;$\mathrm{dB}\text{/}$$ but at that value the estimation of the distance is imprecise and the excess noise too high to get a secret key. However, the distance could be improved either by improving the DSP or by changing the generation of the pilots. Indeed, we believe that one of the limitations comes from the finite amplitude difference that can be applied between the quantum symbols and the pilots. A solution could be to generate the pilots and the quantum symbols on different paths, or to send the pilots on the other polarisation, which would allow for more powerful pilots and less crosstalk, as suggested in [39].
@@
The choice of the optimised parameters reported in this work was performed using a slightly older system with a DAC that was limited to $500\;$\mathrm{MSa}\text{/}\mathrm{s}$$ (Keysight M330A) and with a bandwidth of $200\;$\mathrm{MHz}\text{/}$$. This justifies the relatively low symbol rate used in the present work. The increase of the symbol rate (and hence the occupation bandwidth), allowed by the new DAC will be the subject of future investigation. We believe that there are no particular limitations coming from our software other than longer time and larger memory occupation for the DSP.
@@
Currently, the DSP (and the parameter estimation) is run for each frame before requesting the next frame, which means that two frames are temporally separated by the DSP time. We conducted an experiment on 20 frames to evaluate the DSP time and we measured two durations: the first is the DSP itself, which outputs the symbols for each subframe with a global phase difference, and the second is recovery of the global phase, which also requires Alice sending her data to Bob. The experiment is conducted with the same parameters as the experiment for $0\;$\mathrm{km}\text{/}$$ and requests $50\%$ of the symbols to Alice ($500000$ symbols per frame). The average time for the DSP is $70\;$\mathrm{s}\text{/}\mathrm{frame}$$ and the average time for the symbol exchange and global phase correction is $164\;$\mathrm{s}\text{/}\mathrm{frame}$$ for a total time between frames of $234\;$\mathrm{s}\text{/}\mathrm{frame}$$. The second time can be reduced by exchanging less data for the parameter estimation. For example, taking $10\%$ of the data (i.e. $100000$ symbols for the parameter estimation), the DSP time is similar ($72\;$\mathrm{s}\text{/}\mathrm{frame}$$ on average) and the data exchange and global phase recovery has an average time of $101\;$\mathrm{s}\text{/}\mathrm{frame}$$, yielding a total of $173\;$\mathrm{s}\text{/}$$ between the frames. This has the advantage of reducing the time and increasing the key rate, but the estimators for the finite-size analysis are worse. Both times could be reduced by optimizing the code, or by using a different programming language.
6 Conclusion
We have presented a highly modular open source platform for experimental continuous-variable quantum key distribution that is able to exchange a secret key rate at metropolitan distances using the OSSB-GCMS-GG02 protocol with frequency multiplexed pilots and RF-heterodyne coherent detection. We obtain key rates of the order of the $\mathrm{Mbit}\text{/}\mathrm{s}$ for distances up to $25\;$\mathrm{km}\text{/}$$ at the asymptotic limit. Our platform can be easily extended to other CV-QKD protocols and hardware configurations, including with double polarisation and phase-diverse heterodyne detection. Our platform can be improved and extended by integrating techniques accessible in the near term.